ZYX Inc. shows the following balance sheet as of 12/31/2016 (in US $ millions):

Assets Liabilities
Cash & checking deposits 3 Short-term debt 10
Accounts receivable 4 Accounts payable 6
Inventories & supplies 8 Mortgage debt 14
Vehicles 12 Total debt 30
Machinery & equipment 16
Buildings & structures 24 Equity 37
Total assets 67 Total liabilities 67

ZYX's chief financial officer anticipates that the income statement from 1/1/2017 through 12/31/2017 will show the following figures (US $ millions):

(+) Total revenues 50
(-) Cost of goods sold 22
(-) Operating expenses (incl. depreciation) 10
(-) Interest payments & other financial expenses 4
(-) Taxes 2

If ZYX's chief financial officer's expectations are fulfilled, what will the value of the company's total assets be as of 12/31/2017. Assume that total debt (the sum of short-term, accounts payable, and mortgages) stays at the same level as on 12/31/2016. (The possible answers are all in millions of US $.)
Answer #1

Balance sheet

Assets Amount Liabilities Amount

Cash & checking deposits

3

Accounts receivable

4

Short-term debt

10

Inventories & supplies

8

Accounts payable

6

Fixed assets

Mortgage debt

14

new FA asset

22

Total debt

30

Vehicles , Machinery & equipment &

Buildings & structures

42

Equity 37

Reserve & surplus

12

Total assets 79

Total liabilities

79


* Total assets = 12+16+24 = 52

total assets - dep = 52-10= 42

Therefore value of assets = $79 millions we can also adjust difference of balance sheet in cash or debtors after taking suitable assumptions

Revenue Account

(+) Total revenues

50

(-) Cost of goods sold

22

(-) Operating expenses (incl. depreciation)

10

(-) Interest payments & other financial expenses

4

(-) Taxes

2

Profit 12

It is assumed that operating Expenses includes only Depreciation expenses.